OpenStudy (anonymous):

A hybrid car with a 9.80 gal tank consumes gasoline at a rate of 54.1 miles/gal. How many liters of gasoline will be consumed traveling 132 km? Is my answer correct? Will show work in a post.

OpenStudy (anonymous):

1 mile = 5280 ft 1 ft = 12 in 1 in = 2.54 cm 100 cm = 1 m 1000 m = 1 km \[(\frac{ 54.1 miles }{ gal })(\frac{ 5280 ft }{ mile })(\frac{ 12 inches }{ ft })(\frac{ 2.54cm }{ inch })(\frac{ meter }{ 100cm })(\frac{ 132km }{ 1000m })(\frac{ 1gal }{ 3.78541 L})\] = 3036.037674 since 54.1 is 3 significant figures =3040L I got this wrong (on an online quiz) can you please tell me why. Thanks in advance

OpenStudy (anonymous):